# The Picard group of the Baily-Borel compactification of the moduli space of polarized K3 surfaces

### v1

@@ 2.1 Weil representation
Let $G$ be a discriminant form. For each $d\in\mathbb{Z}_{\geq 1}$, there is a unitary representation $\rho_{G}^{(d)}$ of ${\rm Mp}_{2d}(\mathbb{Z})$ on the group ring $\mathbb{C}[G^{(d)}]$, called the Weil representation associated to $G$. It is defined by
$\rho_{G}^{(d)}(n(B))\mathfrak{e}_{\gamma}$ $=\mathbf{e}\left(\frac{1}{2}\mathrm{tr}((\gamma,\gamma)B)\right)\mathfrak{e}_{\gamma},$ (2.1)
$\rho_{G}^{(d)}(J_{d})\mathfrak{e}_{\gamma}$ $=\frac{\mathbf{e}\left(-\frac{d}{8}\mathrm{sign}(G)\right)}{|G|^{\frac{d}{2}}}\sum\limits_{\delta\in G^{(d)}}\mathbf{e}(-\mathrm{tr}(\gamma,\delta))\mathfrak{e}_{\delta},$
where $\mathfrak{e}_{\gamma}$ is the standard basis of $\mathbb{C}[G^{(d)}]$ for $\gamma=(\gamma_{1},\ldots,\gamma_{d})\in G^{(d)}$.
@@
For $m(U)=\Big{(}\left(\begin{array}[]{cc}U&0\\ 0&(U^{-1})^{\top}\end{array}\right),\sqrt{\det U^{-1}}\Big{)}\in{\rm Mp}_{2n}(\mathbb{Z})$, we have
@@
  • When $G\cong M^{\vee}/M$ for some lattice $M$, we may simply write it as $\rho_{M}^{(d)}$. And if furthermore $d=1$, we may simplify it as $\rho_{M}$.
2.2 Siegel modular forms
For any $k\in\frac{1}{2}\mathbb{Z}$, define the Petersson slash operator $\mid_{k,G^{(d)}}$ on the space of vector-valued functions $f:\mathbb{H}_{d}\rightarrow\mathbb{C}[G^{(d)}]$ by
$(f\mid_{k,G^{(d)}}[\tilde{g}])(\tau):=\phi_{g}(\tau)^{-2k}\cdot\det(g)^{k/2}\cdot\rho_{G}^{(d)}(\tilde{g})^{-1}f(g\tau),$
with $\tilde{g}=(g,\phi_{g}(\tau))\in\widetilde{{\rm GL}_{2}^{+}}(\mathbb{R})$ where
$\widetilde{{\rm GL}_{2}^{+}}(\mathbb{R}):=\left\{\Big{(}g=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in{\rm GL}_{2}^{+}(\mathbb{R}),\phi_{g}(\tau)\Big{)}\mid\exists t\in\mathbb{C}^{\times},\phi_{g}(\tau)^{2}=t\cdot(c\tau+d)\right\}$
is the metaplectic $\mathbb{C}^{\times}$-extension of ${\rm GL}_{2}^{+}(\mathbb{R})$  [^1] This is different from the definition given by Shimura in [31]. Here we adopt the convention of Bruinier-Stein in [11]..
@@ 2.2 Siegel modular forms
A Siegel modular form $f(\tau)$ of weight $k$ and type $\rho_{G}^{(d)}$ is a vector-valued function
$f:\mathbb{H}_{d}\rightarrow\mathbb{C}[G^{(d)}]$
satisfying the following two conditions.
  • For all $\tilde{g}=(g,\phi_{g}(\tau))\in{\rm Mp}_{2d}(\mathbb{Z})$, we have $f\mid_{k,G^{(d)}}[\tilde{g}]=f$.
  • The function $f$ is holomorphic on $\mathbb{H}_{d}$ and at $\infty$.
When $G\cong G_{M}$ for an even lattice $M$, a Siegel modular form of type $\rho_{G}^{(d)}$ has a Fourier expansion
$f(\tau)=\sum\limits_{\gamma\in G^{(d)}}~{}\sum_{\begin{subarray}{c}T\in{\rm Sym}_{d}(\mathbb{Z})_{\rm{even}}+(\gamma,\gamma),\\ T\geq 0\end{subarray}}c(T,\gamma)\mathbf{e}\left(\frac{1}{2}\mathrm{tr}(T\tau)\right)\mathfrak{e}_{\gamma}.$
It is called a cusp form if $c(T,\gamma)\neq 0$ implies $T>0$.
We set ${\mathrm{Mod}}_{k}(\rho_{G}^{(d)})$ (resp. ${\mathrm{Cusp}}_{k}(\rho_{G}^{(d)})$) to be the space of Siegel modular forms (resp. cusp forms) of weight $k$ and type $\rho_{G}^{(d)}$. It is a standard fact that ${\mathrm{Mod}}_{k}(\rho_{G}^{(d)})=0$ if $2k+\mathrm{sign}(G)\nequiv 0\mod 2$. To see this, note that the element $(\mathrm{I}_{2d},-1)$ acts on $\mathbb{C}[G^{(d)}]$ by a scalar $(-1)^{\mathrm{sign}(G)}$. From now on, we assume that the parity condition
$2k+\mathrm{sign}(G)\equiv 0\mod 2$ (2.2)
always holds.
When $d=1$, let
$\langle f,g\rangle_{\rm Pet}:=\int_{{\rm SL}_{2}(\mathbb{Z})\backslash\mathbb{H}}\langle f(z),g(z)\rangle y^{k}\frac{\mathrm{d}x\mathrm{d}y}{y^{2}}$
be the Petersson inner product on ${\mathrm{Mod}}_{k}(\rho_{M})$, which is well-defined if at least one of $f,g$ is a cusp form.
2.3 Some functoriality results regarding Weil representations
@@
    $\downarrow_{H}(\mathfrak{e}_{\gamma})=\mathfrak{e}_{\gamma+H}$
    if $\gamma\in H^{\perp}$ and 00 otherwise.
It is well-known that they are adjoint with respect to the inner product on $\mathbb{C}[G]$ and $\mathbb{C}[H^{\perp}/H]$. Further, they commute with Weil representations, which indicates that they take modular forms to modular forms and preserve cuspidality.
@@
$\Big{(}\left(\begin{matrix}A&B\\ C&D\end{matrix}\right),\phi\Big{)},\Big{(}\left(\begin{matrix}A^{\prime}&B^{\prime}\\ C^{\prime}&D^{\prime}\end{matrix}\right),\phi^{\prime}\Big{)}\longmapsto\Bigg{(}\left(\begin{matrix}A&&B&\\ &A^{\prime}&&B^{\prime}\\ C&&D&\\ &C^{\prime}&&D^{\prime}\end{matrix}\right),\tilde{\phi}\Bigg{)},$
where $\tilde{\phi}(\mathtt{diag}(z,z^{\prime}))=\phi(z)\cdot\phi^{\prime}(z^{\prime})$. They satisfy
$\rho_{M}^{(d)}(\iota(\tilde{g}_{1},\tilde{g}_{2}))(\mathfrak{e}_{\gamma}\otimes\mathfrak{e}_{\gamma^{\prime}})=\rho_{M}^{(r)}(\tilde{g}_{1})(\mathfrak{e}_{\gamma})\otimes\rho_{M}^{(d-r)}(\tilde{g}_{2})(\mathfrak{e}_{\gamma^{\prime}}).$
(cf. [19, Lemma 2.2])
2.4 Siegel-Eisenstein series and Theta series
Suppose $M$ is a negative definite even lattice of rank $r$ under the pairing (,)(-,-)( - , - ). Embedd $M$ into $\mathbb{R}^{r}$ so that the pairing (,)(-,-)( - , - ) is the standard scalar product on $\mathbb{R}^{r}$.
Siegel-Eisenstein series
@@ 2.4 Siegel-Eisenstein series and Theta series
$\left(\begin{array}[]{cc}\rm{I}_{d}&B\\ 0&\rm{I}_{d}\end{array}\right)$ and $\left(\begin{array}[]{cc}U&0\\ 0&(U^{-1})^{\top}\end{array}\right)$
for $B\in{\rm Sym}_{d}(\mathbb{Z})$ and $U\in{\rm GL}_{d}(\mathbb{Z})$. Let $\widetilde{\Gamma}_{\infty}^{(d)}\subseteq{\rm Mp}_{2d}(\mathbb{Z})$ be the preimage of $\Gamma_{\infty}^{(d)}$. For $k\in\frac{1}{2}\mathbb{Z}$ and $k>d+1$, the vector-valued Siegel-Eisenstein series is defined by
@@
Let $H(X)$ be a harmonic polynomial in the matrix variable $X=(X_{ij})_{\begin{subarray}{c}i=1,\dots,r\\ j=1,\dots,d\end{subarray}}$ of degree $h$. We define the Theta series
$\Theta_{M,H}^{(d)}(\tau)=\sum\limits_{\mathbf{v}\in(M^{\vee})^{(d)}}H(\mathbf{v})\mathbf{e}\left(-\frac{1}{2}\mathrm{tr}((\mathbf{v},\mathbf{v})\tau)\right)\mathfrak{e}_{\mathbf{v}+M^{d}}$ (2.4)
@@
$\Theta_{\mathbf{Gen}(M)}^{(d)}=\frac{\sum\limits_{L\in\mathbf{Gen}(M)}|\mathrm{Aut}(L)|^{-1}\sum\limits_{\sigma\in\mathrm{Iso}(G_{M},G_{L})}|\mathrm{Aut}(G_{L})|^{-1}\sigma^{*}\Theta^{(d)}_{L,1}}{|\mathrm{Iso}(G_{M},G_{M})|\sum\limits_{L\in\mathbf{Gen}(M)}|\mathrm{Aut}(L)|^{-1}},$ (2.5)
where $\mathbf{Gen}(M)$ is short for $\mathbf{Gen}_{0,r}(G_{M})$ and $\sigma^{*}\mathfrak{e}_{\gamma}=\mathfrak{e}_{\sigma^{-1}(\gamma)}$ for $\gamma\in G_{L}^{(d)}$.
By [5, Theorem 4.1], $\Theta^{(d)}_{M,H}$ is an element in $\mathrm{Mod}_{h+\frac{r}{2}}(\rho_{M}^{(d)})$. Moreover, $\Theta^{(d)}_{M,H}$ is a cusp form if $h>0$. We define ${\mathrm{Mod}}^{\theta}_{k}(\rho_{M}^{(d)})\subseteq{\mathrm{Mod}}_{k}(\rho_{M}^{(d)})$ to be the subspace
$\mathrm{Span}\left\{\sigma^{*}\Theta_{L,H}^{(d)}\mid~{}L\in\mathbf{Gen}(M),H\hbox{ harmonic of degree }k-\frac{r}{2},\sigma\in\mathrm{Iso}(G_{M},G_{N})\right\},$
and define ${\mathrm{Cusp}}^{\theta}_{k}(\rho_{M}^{(d)})$ as the subspace ${\mathrm{Mod}}^{\theta}_{k}(\rho_{M}^{(d)})\cap{\mathrm{Cusp}}_{k}(\rho_{M}^{(d)})$. The main question we will consider in the following two sections is whether the equality ${\mathrm{Cusp}}^{\theta}_{k}(\rho_{M}^{(d)})={\mathrm{Cusp}}_{k}(\rho_{M}^{(d)})$ holds. In other words, we will consider whether any cusp form can be written as a linear combination of Theta series. For lattices of even rank this was done in [25]. We will adapt the proof to the odd rank case.
@@
When $d=1$, the space of harmonic polynomials in $r$ variables of degree $h$ has dimension $\binom{r+h-1}{r-1}-\binom{r+h-3}{r-1}$ and there is an explicit basis via the Kelvin transform (cf. [2, Theorem 5.25]). When $h=2$, the harmonic polynomials are linear combinations of
$H_{u}(v)=\langle u,v\rangle^{2}-\frac{\langle u,u\rangle\langle v,v\rangle}{r}$ (2.6)
@@
One can identify the genus Theta series with the Siegel-Eisenstein series using the Siegel-Weil formula.
@@
Suppose $M$ is a negative definite even lattice of rank $r$ with $\frac{r}{2}>d+1$. Then
@@
The proof will be given in Appendix A. See Theorem A.5. ∎
3 Vector-valued Hecke operator
3.1 Bruinier-Stein’s vector-valued Hecke operator
Let $G$ be a finite abelian group equipped with a discriminant form whose level is $\ell(G)=:N$. Bruinier and Stein have introduced a Hecke operator acting on ${\mathrm{Mod}}_{k}(\rho_{G})$. For each integer $\alpha>0$, set

### v2

@@ 2.1. Weil Representation
Let $G$ be a discriminant form. For each $d\in\mathbb{Z}_{\geq 1}$, the Weil representation $\rho_{G}^{(d)}$ is a unitary representation of ${\rm Mp}_{2d}(\mathbb{Z})$ on the group ring $\mathbb{C}[G^{(d)}]$. It is defined by
$\rho_{G}^{(d)}(n(B))\mathfrak{e}_{\gamma}$ $=\mathbf{e}\left(\tfrac{1}{2}\mathrm{tr}\left((\gamma,\gamma)B\right)\right)\mathfrak{e}_{\gamma},$ (2.1)
$\rho_{G}^{(d)}(J_{d})\mathfrak{e}_{\gamma}$ $=\frac{\mathbf{e}\left(-\frac{d}{8}\mathrm{sign}(G)\right)}{|G|^{d/2}}\sum_{\delta\in G^{(d)}}\mathbf{e}\bigl{(}-\mathrm{tr}(\gamma,\delta)\bigr{)}\mathfrak{e}_{\delta},$
where $\mathfrak{e}_{\gamma}$ denotes the standard basis vector in $\mathbb{C}[G^{(d)}]$ for $\gamma=(\gamma_{1},\dots,\gamma_{d})\in G^{(d)}$.
@@
For $m(U)=\Bigl{(}\left(\begin{array}[]{cc}U&0\\ 0&(U^{-1})^{\top}\end{array}\right),\sqrt{\det U^{-1}}\Bigr{)}\in{\rm Mp}_{2d}(\mathbb{Z})$, we have
@@
When $G\cong M^{\vee}/M$ for a lattice $M$, we write $\rho_{M}^{(d)}$ for $\rho_{G}^{(d)}$. For $d=1$, we simplify this to $\rho_{M}$.
2.2. Siegel Modular Forms
For $k\in\frac{1}{2}\mathbb{Z}$, define the Petersson slash operator $\mid_{k,G^{(d)}}$ on vector-valued functions $f\colon\mathbb{H}_{d}\to\mathbb{C}[G^{(d)}]$ by
$(f\mid_{k,G^{(d)}}[\widetilde{g}])(\tau)\coloneqq\phi_{g}(\tau)^{-2k}\det(g)^{k/2}\rho_{G}^{(d)}(\widetilde{g})^{-1}f(g\tau),$
where $\widetilde{g}=(g,\phi_{g}(\tau))\in\widetilde{{\rm GL}_{2}^{+}}(\mathbb{R})$ with
$\widetilde{{\rm GL}_{2}^{+}}(\mathbb{R})\coloneqq\Biggl{\{}\Bigl{(}g=\begin{pmatrix}a&b\\ c&d\end{pmatrix},\phi_{g}(\tau)\Bigr{)}\in{\rm GL}_{2}^{+}(\mathbb{R})\times\mathcal{O}(\mathbb{H}_{d})\Bigm{|}\exists t\in\mathbb{C}^{\times},\,\phi_{g}(\tau)^{2}=t(c\tau+d)\Biggr{\}}$
denoting the metaplectic $\mathbb{C}^{\times}$-extension of ${\rm GL}_{2}^{+}(\mathbb{R})$ [^1] This convention differs from Shimura’s [Shimura73], following Bruinier–Stein [BS]..
@@ 2.2. Siegel Modular Forms
A Siegel modular form of weight $k$ and type $\rho_{G}^{(d)}$ is a holomorphic function $f\colon\mathbb{H}_{d}\to\mathbb{C}[G^{(d)}]$ satisfying:
  1. (i)
    $f\mid_{k,G^{(d)}}[\widetilde{g}]=f$ for all $\widetilde{g}\in{\rm Mp}_{2d}(\mathbb{Z})$;
  2. (ii)
    $f$ is holomorphic at cusps.
Such forms admit Fourier expansions
$f(\tau)=\sum_{\gamma\in G^{(d)}}\sum_{\begin{subarray}{c}T\in{\rm Sym}_{d}(\mathbb{Z})_{\mathrm{even}}+(\gamma,\gamma)\\ T\geq 0\end{subarray}}c(T,\gamma)\mathbf{e}\bigl{(}\tfrac{1}{2}\mathrm{tr}(T\tau)\bigr{)}\mathfrak{e}_{\gamma}.$
We call $f$ a cusp form when $c(T,\gamma)\neq 0\Rightarrow T>0$.
Denote by ${\mathrm{Mod}}_{k}(\rho_{G}^{(d)})$ and ${\mathrm{Cusp}}_{k}(\rho_{G}^{(d)})$ the spaces of Siegel modular forms and cusp forms, respectively. The central element $(\mathrm{I}_{2d},-1)$ acts via $(-1)^{\mathrm{sign}(G)}$, implying
${\mathrm{Mod}}_{k}(\rho_{G}^{(d)})=0\quad\text{if}\quad 2k+\mathrm{sign}(G)\not\equiv 0\pmod{2}.$
We therefore universally assume the parity condition
$2k+\mathrm{sign}(G)\equiv 0\pmod{2}.$ (2.2)
For $d=1$, the Petersson inner product on ${\mathrm{Mod}}_{k}(\rho_{M})$ is
$\langle f,g\rangle_{\mathrm{Pet}}\coloneqq\int_{{\rm SL}_{2}(\mathbb{Z})\backslash\mathbb{H}}\langle f(z),g(z)\rangle y^{k-2}\mathrm{d}x\mathrm{d}y,$
well-defined when at least one of them is cuspidal.
2.3. Some functoriality results regarding Weil representations
@@
    $\downarrow_{H}(\mathfrak{e}_{\gamma})=\begin{cases}\mathfrak{e}_{\gamma+H}&\text{if }\gamma\in H^{\perp},\\ 0&\text{otherwise}.\end{cases}$
It is well known that they are adjoint with respect to the standard inner products on $\mathbb{C}[G]$ and $\mathbb{C}[H^{\perp}/H]$ respectively. Further, they commute with Weil representations, which implies that they take modular forms to modular forms and preserve cuspidality.
@@
$\Big{(}\left(\begin{matrix}A&B\\ C&D\end{matrix}\right),\phi\Big{)},\Big{(}\left(\begin{matrix}A^{\prime}&B^{\prime}\\ C^{\prime}&D^{\prime}\end{matrix}\right),\phi^{\prime}\Big{)}\longmapsto\Bigg{(}\left(\begin{matrix}A&&B&\\ &A^{\prime}&&B^{\prime}\\ C&&D&\\ &C^{\prime}&&D^{\prime}\end{matrix}\right),\widetilde{\phi}\Bigg{)},$
where $\widetilde{\phi}(\mathtt{diag}(z,z^{\prime}))=\phi(z)\cdot\phi^{\prime}(z^{\prime})$. This embedding satisfies the following compatibility with the Weil representation
$\rho_{G}^{(d)}(\iota(\tilde{g}_{1},\tilde{g}_{2}))(\mathfrak{e}_{\gamma}\otimes\mathfrak{e}_{\gamma^{\prime}})=\rho_{G}^{(r)}(\tilde{g}_{1})(\mathfrak{e}_{\gamma})\otimes\rho_{G}^{(d-r)}(\tilde{g}_{2})(\mathfrak{e}_{\gamma^{\prime}}),$
where we identify $\mathbb{C}[G^{(r)}]\otimes\mathbb{C}[G^{(d-r)}]$ with $\mathbb{C}[G^{(d)}]$.(cf. [BZ24, Lem. 2.2])
2.4. Siegel–Eisenstein series and theta series
Suppose $M$ is a positive definite even lattice of rank $r$ with bilinear form (,)(-,-)( - , - ). Fix an embedding of $M$ into $\mathbb{R}^{r}$ such that (,)(-,-)( - , - ) coincides with the standard inner product on $\mathbb{R}^{r}$.
Siegel–Eisenstein series
@@ 2.4. Siegel–Eisenstein series and theta series
$\left(\begin{array}[]{cc}\mathrm{I}_{d}&B\\ 0&\mathrm{I}_{d}\end{array}\right)\text{ and }\left(\begin{array}[]{cc}U&0\\ 0&(U^{-1})^{\top}\end{array}\right)$
for $B\in{\rm Sym}_{d}(\mathbb{Z})$ and $U\in{\rm GL}_{d}(\mathbb{Z})$. Let $\widetilde{\Gamma}_{\infty}^{(d)}\subseteq{\rm Mp}_{2d}(\mathbb{Z})$ be the preimage of $\Gamma_{\infty}^{(d)}$ under the metaplectic cover. For $k\in\frac{1}{2}\mathbb{Z}$ and $k>d+1$, the vector-valued Siegel–Eisenstein series is defined by
@@
Let $F(X)$ be a harmonic polynomial in the matrix variable $X=(X_{ij})_{\begin{subarray}{c}i=1,\dots,r\\ j=1,\dots,d\end{subarray}}$ of degree $h$. We define the theta series
$\Theta_{M,F}^{(d)}(\tau)=\sum_{\mathbf{v}\in(M^{\vee})^{(d)}}F(\mathbf{v})\mathbf{e}\left(\frac{1}{2}\mathrm{tr}\left((\mathbf{v},\mathbf{v})\tau\right)\right)\mathfrak{e}_{\mathbf{v}+M^{d}}$ (2.4)
@@
$\Theta_{\mathbf{Gen}(M)}^{(d)}=\frac{\sum\limits_{L\in\mathbf{Gen}(M)}|\mathrm{Aut}(L)|^{-1}\sum\limits_{\sigma\in\mathrm{Iso}(G_{M},G_{L})}\sigma^{*}\Theta^{(d)}_{L,1}}{|\mathrm{Aut}(G_{M})|\sum\limits_{L\in\mathbf{Gen}(M)}|\mathrm{Aut}(L)|^{-1}},$ (2.5)
where $\mathbf{Gen}(M)$ denotes $\mathbf{Gen}_{r,0}(G_{M})$, and $\sigma^{*}\mathfrak{e}_{\gamma}=\mathfrak{e}_{\sigma^{-1}(\gamma)}$ for $\gamma\in G_{L}^{(d)}$.
By [Bo98, Thm. 4.1], $\Theta^{(d)}_{M,F}\in\mathrm{Mod}_{h+\frac{r}{2}}(\rho_{M}^{(d)})$. Moreover, $\Theta^{(d)}_{M,F}$ is a cusp form when $h>0$. Define the subspaces
${\mathrm{Mod}}^{\theta}_{k}(\rho_{M}^{(d)})\coloneqq\mathrm{Span}\left\{\sigma^{*}\Theta_{L,F}^{(d)}\;\middle|\;\begin{array}[]{c}L\in\mathbf{Gen}(M),\\ F\text{ harmonic of degree }k-\frac{r}{2},\\ \sigma\in\mathrm{Iso}(G_{M},G_{L})\end{array}\right\}\subseteq{\mathrm{Mod}}_{k}(\rho_{M}^{(d)}),$
and ${\mathrm{Cusp}}^{\theta}_{k}(\rho_{M}^{(d)})\coloneqq{\mathrm{Mod}}^{\theta}_{k}(\rho_{M}^{(d)})\cap{\mathrm{Cusp}}_{k}(\rho_{M}^{(d)})$. The central question addressed in the next two sections is whether
${\mathrm{Cusp}}^{\theta}_{k}(\rho_{M}^{(d)})={\mathrm{Cusp}}_{k}(\rho_{M}^{(d)}),$
i.e., whether every cusp form is a linear combination of theta series. This was established for even-rank lattices in [Ma24]. We adapt the proof to the odd-rank case.
For a negative definite lattice $M$, define
$\Theta_{M,F}^{(d)}\coloneqq\Theta_{M(-1),F}^{(d)}\in\mathrm{Mod}_{h+\frac{r}{2}}(\rho_{M(-1)}^{(d)})=\mathrm{Mod}_{h+\frac{r}{2}}(\rho_{M}^{(d)*}),$
where $\rho_{M}^{(d)*}$ denotes the dual representation of $\rho_{M}^{(d)}$.
@@
When $d=1$, the space of harmonic polynomials in $r$ variables of degree $h$ has dimension
$\binom{r+h-1}{r-1}-\binom{r+h-3}{r-1}.$
An explicit basis can be constructed via the Kelvin transform (cf. [ABR01, Thm. 5.25]). For $h=2$, harmonic polynomials are linear combinations of
$F_{u}(v)=\langle u,v\rangle^{2}-\frac{\langle u,u\rangle\langle v,v\rangle}{r}$ (2.6)
@@
The genus theta series can be identified with the Siegel–Eisenstein series via the Siegel–Weil formula.
@@
Let $M$ be a positive definite even lattice of rank $r$ with $\frac{r}{2}>d+1$. Then
@@
The proof appears in Appendix A (Theorem A.5). ∎
3. Vector-valued Hecke operator
3.1. Bruinier–Stein’s vector-valued Hecke operator
Let $G$ be a finite abelian group equipped with a discriminant form of level $\mathfrak{n}(G)=:N$. Bruinier and Stein have introduced a Hecke operator acting on ${\mathrm{Mod}}_{k}(\rho_{G})$. For each positive integer $\alpha>0$, set